### v3

@@ 3.2 Key Enabling Technologies
This section describes the key enabling technologies for NET4AI. Note that each of the five capability/service planes in future 6G network architecture as presented in Section 3.1 shall contain its own enabling technologies as we compile in Figure 15. In the following, we only focus on the ones that are closely related to the needs of AI, introducing their research background, research significance, and research status.
@@
3.2.1 Distributed Intelligence and Federated Learning
The NET4AI architecture must address the dual demands of low-latency, high-reliability communication and computing, as well as growing concerns about privacy. Distributed intelligence and FL are the key technologies that can solve these challenges. conventional centralized architectures in wireless networks can no longer meet low-latency, high-reliability communication and computation demands. Additionally, conventional centralized architectures are inefficient in supporting the ubiquitous intelligence required by future wireless networks. This trend towards deploying intelligent decentralized elements has become apparent, with intelligence gradually shifting from centralized to distributed systems. Therefore, introducing distributed intelligent computing architectures is necessary to fully utilize the multi-dimensional data and computing resources held by user terminals and nodes, enabling intelligent connectivity.
@@ 3.2.1 Distributed Intelligence and Federated Learning
Distributed intelligence is a key technology for achieving the intrinsic intelligence of 6G networks. It adopts a decentralized architecture, dispersing network functions and resources across multiple nodes or devices, eliminating the dependence on a single central node to control the entire network. Nodes in a distributed network possess a certain degree of autonomy, enabling them to make decisions and execute tasks independently without frequent communication or coordination with other nodes. This autonomy enhances network flexibility, allowing it to better adapt to various environmental and demand changes. Data is typically stored across multiple nodes in a distributed network, and computational tasks can be distributed among these nodes for processing. This distributed data storage and processing method improves network scalability, accelerates the training and inference processes of large-scale models, and enhances the efficiency of AI systems. Distributed intelligence also supports online learning and continuous optimization by collecting and processing data in real-time, continuously updating models, and improving AI system performance and accuracy. By leveraging the potential of large-scale distributed edge devices, distributed AI technology can overcome the bottlenecks of centralized AI, reduce communication costs, and address data privacy concerns. Distributed AI will build a new intelligent ecosystem for 6G. As shown in Figure 16, the wireless distributed intelligence architectures widely recognized in academia and industry include:
Figure 16: Federated learning, Multi-Agent RL and Split Learning
@@
Federated Learning: FL is a typical paradigm of distributed intelligence, which enhances privacy by training AI models on local devices and only sharing model parameters instead of raw data. During training, each node trains the model locally and uploads the model weights regularly. The central node integrates these weights to form the weights of the federated model and then feeds the weights back to each node for the next round of training or local inference. FL ensures that user data remains local and only transmits weights/gradients to achieve global training, thus enhancing privacy. A typical example of implementing model training using FL in wireless networks is shown in part a of Figure 16. The entire FL process involves downloading the global model $\omega^{t}$ of the current round $t$ to local devices; Device $k$ trains model based on its own private dataset $D_{k}$ to obtain the local gradient update $\mathbf{g}^{t}_{k}$ and then uploads it; The parameter server executes the federated averaging (FedAvg) algorithm to obtain the gradient update aggregation $G^{t}$ and updates the new global model $\omega^{t+1}$ for round $t+1$. This process will be iterated until the end. In addition, FL can be further divided into two types according to the correspondence between the feature dimensions of the dataset used and the samples: horizontal FL and vertical FL.
@@
(1) Horizontal FL: Also known as sample-based FL, this can be applied in scenarios where the datasets of the various participants in FL have the same feature space but different sample spaces. Through horizontal FL, different participants can jointly train a more powerful AI model based on local data sets while protecting privacy and improving the accuracy and generalization of the model. Additionally, horizontal FL enables real-time learning on distributed edge devices, allowing for timely model parameter updates to constantly changing environments and data. This is crucial for AI applications that require quick responsiveness and strong adaptability. It also supports collaboration across devices with varying computational capabilities and data distributions, catering to the diverse needs of 6G network terminals. Personalized model training on local devices enhances user experience by offering more individualized services.
@@
The authors of [141] took into account the actual situation of limited resources in wireless networks, and deeply analyzed the convergence of FedAvg in horizontal FL under the two assumptions of full and partial device participation, providing important theoretical support for the applicaiton of horizontal FL in wireless networks. In [142, 143], the authors furthermore represented the partial device participation as a probability distribution and eliminate the device selection bias by modifying the aggregation rule in FedAvg to ensure the unbiasedness of gradient updates. This allowed them to provide a more stringent convergence bound and is committed to the efficient implementation of horizontal FL. Based on these theoretical supports, the authors of [144] provided a variety of solutions to guide the efficient application of FL in wireless networks, and was committed to reducing the severe energy consumption problem of FL from aspects such as layering, aggregation frequency optimization, device selection, and communication compression, so as to achieve low-carbon green economic goal. In [145], the authors focued on the shortcoming of horizontal FL in dealing with the heterogeneity of device data, and introduced a personalized federation learning (PFL) method with theoretical performance derivation, providing a theoretical basis for the applicaiton of FL in multiple fields to realize personalized models. The authors of [146] categorized PFL techniques based on key challenges and strategies, highlighting key ideas and future research directions in PFL architecture, realistic benchemarks, and trustworthy methods. In [147], the authors considered the applicaiton of horizontal FL in the medical field to solve the problem of collaborative research and patient health testing among multiple hospitals, and proposed a blockchain-based orchestration ML system framework to explore the protection of medical data privacy. The authors of [148] focued on the field of IoT healthcare and proposed an intelligent medical system agent architecture that integrated horizontal FL and blockchain technology to achieve efficient and secure processing of medical data to optimize the medical diagnosis process. The authors of [149] proposed a new FL framework for cross-domain prediction for intelligent manufacturing, designed to quickly adapt to new products, processes, and applications with limited training data.
Horizontal FL has important application prospects in providing efficient and personalized AI services in wireless networks, coordinating research with multiple hospitals in healthcare, and realizing cross-application in intelligent manufacturing. However, it also has defects such as slow convergence speed and susceptibility to data heterogeneity.
(2) Vertical FL: Vertical FL refers to a setup where datasets have different features but share the same samples/users. It is suitable for scenarios where there is a high overlap of users among participants but low or no overlap in feature space by extracting parts where participants have the same users but not the same features for joint training. In the training process, it is first necessary to sample align each participant’s data to obtain overlapping user data, then train the model locally and upload the parameters to the server [150]. The difference between vertical and horizontal FL is that horizontal FL emphasizes cooperation and sharing between different data sources, while vertical FL focuses on collaboration and sharing within varying levels of the same data source.
Vertical FL can leverage the correlation between different levels of data sources to achieve data augmentation. By sharing model parameters, various data sources can complement and enrich each other’s data, improving AI models’ generalization ability and robustness. Furthermore, vertical FL can facilitate cross-level learning and knowledge transfer. Different data sources can share model parameters and learning experiences, achieving cross-level knowledge transfer and migration and enhancing the overall model’s effectiveness.
The authors of [151] proposed a vertical federated recommendation system method based on clustering and latent factor model. Facing the diversity of a large number of different users, the users are clustered, and the prediction rating model and clustered user latent factor vector are reconstructed to more fully explore the relationship between users and items and improve the accuracy of user recommendations. In [152], the authors described the idea of intent-hidden vertical FL, taking the example of multi-party collaborative training of disease prediction models in medical diagnosis, and buildt a security protocol using homomorphic encryption to meet the exact needs of data privacy protection. The authors of [153] proposed a vertical federated DL based on the spatial-temporal long and short-term networks algorithm for intelligent city traffic in MEC environment. Using different sample features (i.e. different regions) in the same sample space (i.e. the same traffic scene and time range) to build a federated model to predict traffic flow, effectively solving the problems of data privacy and spatio-temporal information capture in traffic prediction, providing more accurate prediction results, and helping to optimize traffic management and improve travel efficiency.
Vertical FL can fully explore the connections between different features of data samples and is of great significance in recommendation systems and traffic prediction. However, it is still in its early stages of research and related theoretical research has not yet been perfected, so it cannot provide sound theoretical support.
Multi-Agent Reinforcement Learning: This branch of RL studies the behavior of multiple learning agents coexisting in a shared environment. Each agent pursues its reward and acts to advance its interests. The primary goal is to achieve cooperative and win-win strategy learning among multiple distributed nodes, suitable for multi-agent collaboration or competition scenarios. 6G networks must support diverse AI application scenarios, such as smart cities and intelligent transportation. The multi-agent reinforcement learning (MARL), with its flexible learning and decision-making capabilities, can offer personalized and optimized solutions across various applications. By learning the characteristics of their environment and tasks, agents can autonomously optimize and adjust to meet different application needs, enhancing user experience.
The authors of [154] proposed an intelligent and decentralized MARL algorithm based on advantage actor critic to solve large-scale traffic signal control problems, providing an effective solution to the non-stationary problem of congestion relief in intelligent transportation systems. Similarly, for the problem of controlling traffic lights in cities, the authors of [155] proposed a new multi-agent recurrent DDPG algorithm, which adjusts the phase of traffic signals in real time through Internet of Vehicles data, effectively reducing traffic congestion under complex road conditions and partially observable environments. In addition, in the intelligent transportation, the authors of [156] proposed the use of partially observable markov game to formulate connected autonomous driving problems with realistic assumptions, and developed a multi-agent interconnected autonomous driving learning platform, providing a new practical basis for the development of autonomous driving technology. In [157], the authors focused on achieving collaborative computing between multiple layers of low earth orbit satellites. By establishing a sequential model for task processing and obtaining the connection relationship between satellites in different orbits, a real-time offloading decision model was constructed and the use of MARL combined with penalty function method was proposed to optimize the average performance of the system under continuous time scales, providing a new solution for improving the system performance of satellites supporting edge computing under non-uniform loads. In [158], the authors introduced a MARL-based online production scheduling method in smart factories. Through distributed computing architecture and compound reward functions, they effectively proved that the AI-based online solution is superior to the conventional solution in production scheduling performance, and shows stronger adaptability and robustness under uncertain situations such as order changes and machine failures. The authors of [159] proposed a novel multi-agent DRL scheme to solve the multi-channel access and task offloading problems in MEC-enabled Industry 4.0 and improved the computational efficiency and resource utilization of industrial systems.
MARL is suitable for signal control and autonomous driving applications in intelligent transportation, and is expected to become a promising solution in collaborative computing between satellites and production scheduling in intelligent industry, resource utilization optimization of task offloading, etc. However, it still has shortcomings such as high algorithm complexity and lack of data privacy protection.
Split Learning: Considering that devices with limited resources cannot support complex DL models and thus cannot support collaboration in the form of FL [160], the industry has derived split learning (SL) technology. SL splits the ML model and allows different parts of the model to be distributed between the client and the server, while allowing collaboration for joint model training without leaking the original data [161]. High computing tasks are assigned to powerful central nodes, while shallower data processing layers are retained at the terminal where the data is stored. This can make ML easier to use for resource-constrained devices. As an alternative/complementary method to FL, SL can provide better privacy protection and reduce the processing load of resource-constrained devices, and has the potential to become one of the key AI technologies for 6G distributed intelligence.
In the field of healthcare, the development of DL is constrained by the lack of data samples and data privacy protection. Based on this, the authors of [162, 163] introduced SL into healthcare applications to solve the problem of different entities collaboratively training DL models without sharing sensitive raw data, and achieved good experimental results on the model performance based on medical datasets. Similarly, in [164], the authors studied the use of SL for multi-institutional collaborative training of DL models, and verified that SL can achieve better client efficiency and more reliable privacy protection. Considering that the sequential training method of original SL may cause model overfitting or slow convergence, the authors of [165] proposed a parallel SL method, which solved the overfitting problem in sequential SL through training batch adaptive adjustment and node layer synchronization mechanism, and improved the training effect while ensuring privacy. In [166], the authors introduced SL into wireless networks to realize the utilization of a large amount of terminal privacy data in IoT. Considering that the sequential training method leads to large delays, they proposed a cluster-based parallel training method, and simultaneously optimized the joint cutting layer selection, device clustering and spectrum allocation to effectively reduce the training delay.
SL is suitable for multi-institutional collaborative training of DL models in healthcare with strict privacy requirements, and is also suitable for training efficient DL models in IoT. However, it has the disadvantages of being easily affected by data heterogeneity, complex training methods, overfitting of models, slow convergence, or even failure to converge.
@@
The above are some typical representative technologies in distributed intelligence. Different technologies have different application scenarios and scopes of application. For FL, it is more suitable for efficiently implementing AI services in scenarios with data privacy requirements. And FL is widely applicable to various emerging intelligent demand applications of 6G owing to its compatibility with other intelligent learning paradigms. For MARL, it is more in line with the needs of online learning in complex environments and completes AI tasks in an observable and interactive environment. For separation learning, it is suitable for realizing collaborative learning of DL models in resource-constrained devices, and can adapt to different levels of resource constraints through multi-level model splitting.
With the development of the networks and the advancement of ML technology, future distributed intelligence may also evolve towards other research directions. (1) To develop more reliable distributed intelligence, more effective error correction and retransmission mechanisms can be adopted to address data loss and errors in unreliable communication environments, such as advanced error correction codes and adaptive retransmission strategies to reduce the error rate and improve communication reliability. (2) To develop more secure distributed intelligence, more secure protocols and privacy protection mechanisms should be further considered. Enhanced data encryption, identity authentication and access control measures can be employed to protect user data from unauthorized access and tampering. In addition, there are some frontier directions that can be explored[125, 167], such as improving the communication efficiency of data/model parameter exchange, enhancing the energy efficiency of learning algorithms, and improving fairness through the incentive mechanisms for participating devices.
@@ 3.2.2 Digital Twins for 6G Network native AI
In terms of management and orchestration services, DT is a key enabling technology that enables more efficient and accurate allocation of resources and system optimization for AI services. In B5G/6G communication systems, the rapid growth in user demand and the continuous evolution of technical standards have posed significant challenges to resource allocation and system optimization. conventional communication systems, which rely on historical data and static models, struggle to adapt swiftly to dynamic network changes, making real-time adjustments difficult. These limitations hinder the adaptability, reliability, and flexibility of communication networks. DT technology has been introduced into the communication domain to address these challenges, offering a flexible, efficient, and precise approach to resource optimization. Adopting DT not only enhances system agility and performance but also introduces innovative models for network operation, significantly reducing research and development costs and risks [168]. Figure 17 depicts the relationship between 6G physical infrastructure and DT for AI.
@@
The concept of the DT was first proposed by Professor Michael Grieves [169]. It represents a technology that leverages sensors, data processing, modeling, and simulation to comprehensively map physical entities’ behavior, performance, and state in a virtual environment. By creating a virtual replica of a physical object, DTs establish a bidirectional connection between the physical and digital worlds, enabling real-time reflection of the physical entity’s operational state and the simulation of its future behavior. The primary objective of DT technology is to enable comprehensive monitoring, prediction, and optimization of physical entities through digitalization and intelligence, thereby enhancing system reliability, flexibility, and performance. In 6G communication systems, DT technology facilitates the construction of comprehensive network digital replicas to monitor network devices in real time, predict network states, and optimize resource allocation [170]. This holistic digital mapping transforms network operations from static planning to dynamic responsiveness.
@@
The implementation of DT technology relies on the synergy of several key technologies, including high-precision data collection, real-time modeling and simulation, bidirectional data interaction, and intelligent analysis and optimization. In communication systems, distributed sensor networks continuously gather diverse data from physical entities, such as network device statuses, traffic dynamics, environmental parameters, and user behavior patterns. This data is transmitted to the DT platform via high-speed, low-latency communication links.
The DT platform employs physical models and data-driven ML algorithms to construct high-fidelity virtual replicas that comprehensively simulate physical entities’ operational state and dynamic changes. These virtual replicas accurately reflect the current state and leverage simulation techniques to predict future system behaviors or potential anomalies. With bidirectional data interaction, physical systems and virtual replicas achieve real-time data synchronization. The DT analyzes large-scale data intelligently, generates optimized decisions, and swiftly relays these to the physical system. For instance, when a network traffic surge is anticipated, the DT can proactively adjust load distribution strategies, optimize spectrum and power resources, and alleviate stress on network devices.
The integration of edge computing further enhances the efficiency of this technology by distributing modeling, simulation, and optimization tasks to edge nodes closer to physical entities. This setup ensures processing speed and responsiveness meet ultra-low latency requirements. Through the seamless integration of these advanced technologies, DT enables intelligent management and dynamic optimization of complex networks in 6G communication systems, providing robust support for efficient resource utilization, network stability, and intelligent upgrades [171].
@@
The 6G network aims to achieve ultra-high capacity, ultra-short-range communication, beyond-best-effort and high-precision communication, and integrate multiple communication types [172]. However, this ambitious vision introduces security, spectrum efficiency, intelligence, energy efficiency, and affordability challenges. The emergence of DT technology offers promising opportunities to address these challenges. DT provides a virtualized counterpart for 6G networks, enabling comprehensive network traffic monitoring and analysis. By collecting traffic data across the entire network, DT utilizes data analytics to proactively uncover traffic patterns and detect anomalies. Leveraging feedback from the virtualized network, 6G systems can enhance their security by preparing for potential threats in advance. Additionally, DT enables the automation of demand identification and service provisioning by analyzing communication data to discern usage patterns and rules. Predictive insights into communication demands allow 6G networks to reserve resources, such as spectrum, in anticipation of future needs. Moreover, the integration of DT technology empowers 6G to support innovative services, including AR/VR, and autonomous driving. By addressing key challenges in security, spectrum efficiency, intelligence, energy efficiency, and customization, DT technology redefines and accelerates the development of 6G networks [173].
@@
Significant research has integrated DT technology with 6G networks to enhance network performance. For instance, DT has been proposed for managing metasurface reflectors in 6G terahertz communications [174]. By modeling, predicting, and controlling the propagation characteristics of indoor signals, DT maximizes the system’s terahertz SNR. Additionally, Lu et al. introduced the concept of DT wireless networks by integrating DT into wireless networks. This approach shifts real-time data processing and computation to the edge plane and leverages DT to mitigate the unreliability of long-distance communications between end users and edge servers in 6G networks [175]. Overall, the deep integration of DT with 6G effectively bridges physical systems and digital spaces, providing robust technical support for efficient and reliable wireless communication.
@@
One of the critical capabilities of 6G networks will be supporting massive-scale terminal devices that generate extensive data traffic. In this context, DT can optimize edge resources to meet user demands better. For example, the authors of [176] designed a DT model for edge computing in 6G wireless networks. By incorporating mobile users’ dynamic network state and DT, the study addressed edge association problems using DRL and transfer learning, successfully reducing average system latency and improving resource utilization. Similarly, the authors of [177] proposed a DT-assisted task offloading method, creating DTs for all device states to achieve lower latency and power consumption in the network. Furthermore, the authors of [178] investigated dynamic resource allocation mechanisms for DT-based services in the 6G IoT. The study improved resource allocation efficiency by establishing service function chains with DTs and employing a collective RL method.

### v4

@@ 3.2 Key enabling technologies
This section describes the key enabling technologies for NET4AI. Each of the five capability/service planes in the future 6G network architecture, as presented in Section 3.1 shall contain its own enabling technologies, as categorized in Figure 13. In the following, we focus on those closely related to the needs of AI, introducing their research background, significance, and progress.
3.2.1 Distributed Intelligence and Federated Learning
The NET4AI architecture must address the dual demands of low latency and high reliability for communication and computing, as well as growing concerns about privacy. Distributed intelligence and FL are the key technologies that can solve these challenges. Conventional centralized architectures in wireless networks can no longer meet low-latency, high-reliability communication and computation demands. Additionally, conventional centralized architectures are inefficient in supporting the ubiquitous intelligence required by future wireless networks. This trend towards deploying intelligent decentralized elements has become apparent, with intelligence gradually shifting from centralized to distributed systems. Therefore, introducing distributed intelligent computing architectures is necessary to fully utilize the multi-dimensional data and computing resources held by user terminals and nodes, enabling intelligent connectivity.
@@ 3.2.1 Distributed Intelligence and Federated Learning
Distributed intelligence is a cornerstone technology for realizing the intrinsic intelligence envisioned for 6G networks. By adopting a decentralized architecture, it disperses network functions and resources across multiple nodes or devices, thereby eliminating reliance on a single central node to manage the entire network. Each node in a distributed network operates with a degree of autonomy, enabling independent decision-making and task execution without the need for constant communication or coordination with other nodes. This autonomy enhances network flexibility, making it more adaptable to dynamic environmental conditions and evolving demand patterns.
In distributed networks, data is stored across multiple nodes, and computational tasks are processed collaboratively among them. This decentralized approach to data storage and computation improves network scalability, accelerates the training and inference of LAMs, and enhances overall system efficiency. Moreover, distributed intelligence facilitates online learning and continuous optimization by enabling real-time data collection, processing, and model updates, which improves the performance and accuracy of AI systems. Distributed AI technology addresses the limitations of centralized AI systems, such as communication bottlenecks and data privacy concerns. It reduces communication overhead and ensures data security by keeping sensitive information closer to its source. Distributed AI is poised to create a new intelligent ecosystem for 6G, fostering a more scalable, efficient, and privacy-preserving network paradigm. As shown in Figure 14, the wireless distributed intelligence architectures include:
@@
(1) Federated Learning. FL is a typical paradigm of distributed intelligence, which enhances privacy by training AI models on local devices and only sharing model parameters instead of raw data [149]. During training, each client node trains the model locally and uploads the model weights regularly. A central node aggregates these weights and feeds the aggregated weights back to the clients for the next round of training or local inference. A typical example of FL in wireless networks is illustrated in Figure 14a. The process begins with downloading the global model $\omega^{t}$ for the current round $t$ to local devices. Each device $k$ trains the model on its private dataset $D_{k}$, generating a local gradient update $\mathbf{g}^{(t)}_{k}$, which is then uploaded to the parameter server. The server applies the federated averaging (FedAvg) algorithm to aggregate these updates into a global gradient $G^{(t)}$, subsequently updating the global model to $\omega^{t+1}$ for the next round. This process continues until model convergence. FL can be divided into horizontal FL and vertical FL. Horizontal FL emphasizes inter-source cooperation and data sharing, whereas vertical FL facilitates collaboration and information exchange across distinct levels within the same data source.
Figure 14: An illustration of examples of FL, MARL, and SL in distributed intelligence.
@@
Horizontal FL. Also known as sample-based FL, horizontal can be applied in scenarios where the datasets of the various participants in FL have the same feature space but different sample spaces. Through horizontal FL, different participants can jointly train a model based on local data sets while protecting privacy and improving the accuracy and generalization of the model. Horizontal FL enables real-time learning on distributed edge devices, allowing for timely model parameter updates to constantly changing environments and data. It supports collaboration across devices with varying computational capabilities and data distributions, catering to the diverse or personalized needs of 6G network terminals [150]. It also supports over-the-air commputations to achieve excellent scalability [151, 152], tolerates to imperfect wireless transmissions [153], and can be potentially implemented in a fully decentralized fashion [154].
@@
The authors of [155] considered limited resources in wireless networks, and analyzed the convergence of FedAvg in horizontal FL under full and partial device participation, providing important theoretical support for the application of horizontal FL in wireless networks. In [156, 157], the authors represented the partial device participation as a probability distribution and eliminated the device selection bias by modifying the aggregation rule in FedAvg to ensure the unbiasedness of gradient updates. This allows for a more stringent convergence bound and is committed to the efficient implementation of horizontal FL. The authors of [158, 159] provided a variety of solutions to guide the efficient application of FL in wireless networks, such as layering, aggregation frequency optimization, device selection, and communication compression, in support of low-carbon green economics. In [160], the authors focused on the shortcomings of horizontal FL in dealing with the heterogeneity of device data, introduced a personalized federated learning (PFL) method, and provided a theoretical basis for the application of FL in multiple fields to realize personalized models. The authors of [161] categorized PFL techniques based on key challenges and strategies, highlighting key ideas and future research directions in PFL architecture, realistic benchmarks, and trustworthy methods. In [162], the authors considered the application of horizontal FL in the medical field to solve the problem of collaborative research and patient health testing among multiple hospitals, and proposed a blockchain-based orchestration framework to enhance medical data privacy. The authors of [163] focused on IoT healthcare and proposed an intelligent medical system agent architecture that integrated horizontal FL and blockchain technology to achieve efficient and secure processing of medical data. The authors of [164] proposed a new FL framework for cross-domain prediction for intelligent manufacturing, designed to quickly adapt to new products, processes, and applications with limited training data. Horizontal FL has important application prospects in providing efficient and personalized AI services in wireless networks, coordinating research with multiple hospitals in healthcare, and realizing cross-application in intelligent manufacturing. However, it also has defects such as slow convergence speed and susceptibility to data heterogeneity and poisoning attacks [165, 166].
Vertical FL. Vertical FL refers to a setup where datasets have different features but the same samples/users. It is suitable for scenarios where there is a high overlap among participants but low or no overlap of their features. During the training process, participant data must first be aligned through sampling to ensure overlapping datasets across users. Subsequently, the model is trained locally, and the resulting parameters are uploaded to a central server [167]. Vertical FL leverages correlations between diverse data sources to achieve data augmentation, enabling complementary and enriched datasets that enhance the generalization and robustness of AI models. By sharing model parameters, vertical FL facilitates cross-level learning and knowledge transfer, allowing data sources to exchange learning experiences and improve overall model effectiveness. To tackle the problem of user diversity, a novel approach for a vertical collaborative recommendation system based on clustering algorithms and latent factor models is put forward [168]. This method effectively enhances the accuracy of the recommendation results. In [169], the authors developed the idea of intent-hidden vertical FL for multi-party collaborative training of disease prediction models in medical diagnosis, and built a security protocol using homomorphic encryption for data privacy protection. The authors of [170] proposed a vertical federated DL based on the spatial-temporal long- and short-term networks for intelligent city traffic in MEC environments. Leveraging diverse sample features (e.g., different regions) within the same sample space (e.g., a specific traffic scene and time range) enables the construction of a federated model for traffic flow prediction. Vertical FL can uncover relationships between different features of data samples, making it highly valuable for applications such as recommendation systems and traffic prediction. Research on vertical FL remains in its infancy, with limited theoretical advancements to date.
(2) Multi-Agent Reinforcement Learning (MARL). In a shared environment, multiple learning agents pursue their own rewards. 6G networks, supporting diverse AI applications like smart cities and intelligent transportation, can benefit from MARL. With its adaptive decision-making, MARL enables agents to autonomously optimize and adjust, enhancing user experience across various applications. The authors of [171] introduced a decentralized MARL algorithm based on advantage actor-critic to address large-scale traffic signal control, effectively tackling congestion in intelligent transportation systems. In [172], a multi-agent recurrent DDPG algorithm was proposed to manage traffic light phases in real-time using Internet of Vehicles data, mitigating congestion in complex and partially observable road environments. In [173], a partially observable Markov game was used to model connected autonomous driving, resulting in a multi-agent learning platform for interconnected autonomous vehicles. The authors of [174] focused on collaborative computing in low earth orbit (LEO) satellites, developing a real-time offloading decision model optimized through MARL with a penalty function. Additionally, an MARL-based online production scheduling method was proposed in [175] for smart factories, while multi-agent DRL was applied in [176] for multi-channel access and task offloading in MEC-enabled Industry 4.0.
(3) Split Learning (SL). Due to the limitations of resource-constrained devices that cannot support complex DL models or FL collaboration [177], SL has emerged as a viable solution [178]. SL divides ML models, distributing different parts between the client and server, ensuring that original data remains secure [179]. While high-computing tasks are handled by powerful central nodes, lighter data processing layers are retained on the terminal where the data are stored. As an alternative or complement to FL, SL reduces the processing burden on resource-limited devices and holds the potential for advancing distributed intelligence in 6G. The authors of [180, 181] introduced SL in healthcare applications, enabling multiple entities to collaboratively train DL models without sharing sensitive raw data, achieving promising performance results on medical datasets. In [182], SL was explored for multi-institutional collaborative training, demonstrating improved client efficiency and more reliable privacy protection. To address issues such as overfitting and slow convergence in the original sequential training method, the authors of [183] proposed a parallel SL method. The method mitigated overfitting through adaptive training batch adjustments and node layer synchronization, improving training efficiency while maintaining privacy. In [184], SL was applied to wireless networks to leverage terminal privacy data in IoT. A cluster-based parallel training method was proposed, optimizing joint cutting layer selection, device clustering, and spectrum allocation to reduce training delays effectively. However, SL remains vulnerable to data heterogeneity, complex training methods, overfitting, slow convergence, and potential failure to converge.
@@
(4) Summary and lessons learned. The above-mentioned technologies can all contribute to distributed intelligence. FL is suitable for AI services in scenarios with data privacy requirements, and applicable to various emerging intelligent demand applications of 6G owing to its compatibility with other intelligent learning paradigms. MARL is in line with the needs of online learning in complex environments and completes AI tasks in an observable and interactive environment. SL is suitable for collaborative learning of DL models in resource-constrained devices, and can adapt to different levels of resource constraints through multi-level model splitting. Distributed intelligence may expand into additional research directions. These include: 1) the adoption of error correction and retransmission mechanisms, such as advanced error correction codes and adaptive retransmission strategies, to ensure reliability in unreliable communication environments; 2) the implementation of secure protocols and privacy protection measures, including data encryption, identity authentication, and access control, to safeguard user data from unauthorized access and tampering; 3) the efficiency of data/model parameter exchange to enhance the energy efficiency of learning algorithms, and 4) fairness through incentive mechanisms for participating devices [185].
@@ 3.2.2 Digital Twins for 6G Network Native AI
Conventional communication systems relying on historical data and static models struggle to adapt swiftly to dynamic network changes, making real-time adjustments difficult. These limitations hinder the adaptability, reliability, and flexibility of the systems. DT technology has been introduced into the communication domain to address these challenges, offering a flexible, efficient, and precise approach to resource optimization. Adopting DT not only enhances system agility but also introduces innovative models for network operation, significantly reducing research and development costs and risks [186]. Figure 15 depicts the relationship between 6G physical infrastructure and DT for AI.
@@
The concept of the DT was first proposed in [187]. By creating a virtual replica of a physical object, DTs establish a bidirectional connection between the physical and digital worlds, enabling real-time reflection of the physical entity’s operational state and the simulation of its future behavior [188]. In 6G communication systems, DT technology facilitates the construction of comprehensive network digital replicas to monitor network devices in real time, predict network states, and optimize resource allocation [189].
@@
The implementation of DT relies on the synergy of several key technologies, including high-precision data collection, real-time modeling and simulation, bidirectional data interaction, and intelligent analysis and optimization. In communication systems, distributed sensors continuously gather diverse data about network device statuses, traffic dynamics, environmental parameters, and user behavior patterns. These data are transmitted to the DT platform via high-speed, low-latency communication links. The DT platform employs physical models and data-driven ML algorithms to construct high-fidelity virtual replicas that comprehensively simulate physical entities’ operational state and dynamic changes [190]. With bidirectional data interaction, physical systems and virtual replicas achieve real-time data synchronization. When a network traffic surge is anticipated, the DT can proactively adjust load distribution strategies, optimize spectrum and power resources, and alleviate stress on network devices. The integration of edge computing further enhances the efficiency of this technology by distributing modeling, simulation, and optimization tasks to edge nodes closer to physical entities. DT enables intelligent management and dynamic optimization of complex networks in 6G communication systems, providing robust support for efficient resource utilization, network stability, and intelligent upgrades [191].
@@
The 6G network aims to achieve ultra-high capacity, ultra-short-range communication, beyond-best-effort service, high-precision communication, and the integration of multiple communication types [192]. However, this vision introduces security, spectrum efficiency, intelligence, energy efficiency, and affordability challenges. The emergence of DT technology offers promising opportunities to address these challenges. DT provides a virtualized counterpart for 6G networks, enabling comprehensive network traffic monitoring and analysis. Leveraging feedback from the virtualized network, 6G systems can enhance their security by preparing for potential threats in advance. Additionally, DT enables the automation of demand identification and service provisioning by analyzing communication data to discern usage patterns and rules. Predictive insights into communication demands allow 6G networks to reserve resources, such as spectrum, in anticipation of future needs. Moreover, the integration of DT technology empowers 6G to support innovative services, including AR/VR, and autonomous driving. By addressing key challenges in security, spectrum efficiency, intelligence, energy efficiency, and customization, DT technology redefines and accelerates the development of 6G networks [193].
@@
Significant research has integrated DT technology with 6G networks to enhance network performance. For instance, DT has been proposed for managing metasurface reflectors in 6G terahertz communications [194]. By modeling, predicting, and controlling the propagation characteristics of indoor signals, DT maximizes the system’s terahertz signal-to-noise ratio (SNR). Additionally, Lu et al. [195] introduced the concept of DT wireless networks by integrating DT into wireless networks. This approach shifts real-time data processing and computation to the edge plane and leverages DT to mitigate the unreliability of long-distance communications between end users and edge servers in 6G networks.
@@
One of the critical capabilities of 6G networks will be supporting massive-scale terminal devices that generate extensive data traffic. For example, the authors of [132] designed a DT model for edge computing in 6G wireless networks. By incorporating mobile users’ dynamic network state and DT, the study addressed edge association problems using DRL and transfer learning, successfully reducing average system latency and improving resource utilization. Similarly, the authors of [196] proposed a DT-assisted task offloading method, creating DTs for all device states to achieve lower latency and power consumption in the network. Furthermore, the authors of [197] investigated dynamic resource allocation mechanisms for DT-based services in the 6G IoT. The study improved resource allocation efficiency by establishing service function chains with DTs and employing a collective RL method.